OpenStudy (khally92):

A bullet of mass 45 g is fired at a speed of 220 m/s into a 5.0-kg sandbag hanging from a string from the ceiling. The sandbag absorbs the bullet and begins to swing. To what maximum vertical height will it rise? Hint: Energy is not conserved as the bullet enters the bag but is conserved after the bullet comes to rest in the bag and the bag is swinging upward

OpenStudy (khally92):

v1'=((m1-m2)/(m1+m2))v1 for the bullet =176ms^-1 And for the sandbag at rest v2'=((2m1)/(m1+m2))v1 =396ms^-1 The bullet is conserved after the bullet comes to rest in the bag. Hence Mgh= 1/2mv^2. I got stuck here what is going to be the value of M, is it sum of the two masses. and value of v as well. And i hope the above is collect. Thanks

OpenStudy (khally92):

OR m1u1+m2u2=(m1+m2)V 0.045(220)+(5)(0)=(5.045)V V=1.96ms^-1 Mgh=1/2mv^2 5.045(9.8)(h)=1/2(5.045)(1.96)^2 h=0.196m Its like this one is more correct. Thanks.

OpenStudy (khally92):

mr Doe and Jonask i await ur reply. Thanks

OpenStudy (anonymous):

2nd one is correctly executed

OpenStudy (khally92):

Thanks bro.